### v1

# Metacognition and Uncertainty Communication in Humans and Large Language Models

Abstract
Metacognition—the capacity to monitor and evaluate one’s own knowledge and performance—is foundational to human decision-making, learning, and communication. As large language models (LLMs) become increasingly embedded in high-stakes decision contexts, it is critical to assess whether, how, and to what extent they exhibit metacognitive abilities. Here, we provide an overview of current knowledge of LLMs’ metacognitive capacities, how they might be studied, and how they relate to our knowledge of metacognition in humans. We show that while humans and LLMs can sometimes appear quite aligned in their metacognitive capacities and behaviors, it is clear many differences remain; attending to these differences is crucial not only for enhancing human-AI collaboration, but also for promoting the development of more capable and trustworthy artificial systems. Finally, we discuss how endowing future LLMs with more sensitive and more calibrated metacognition may also help them develop new capacities such as more efficient learning, self-direction, and curiosity.
1 Introduction
Metacognition refers to the human capacity to monitor, assess, and regulate our own cognitive processes and mental states. Within the broader field of metacognition, confidence judgments—subjective evaluations of the certainty of our own decisions—play a particularly important role (Fleming, 2024). Confidence judgments help individuals assess when their choices are likely correct or potentially erroneous. Expressing confidence also enables effective communication about uncertainty, facilitating collaboration by helping teams leverage collective knowledge and recognize trustworthy decisions (Frith, 2012).
Recent developments in artificial intelligence (AI) have placed considerable attention on uncertainty and its effective communication to human users. Large language models (LLMs), in particular, increasingly serve in advisory roles, providing recommendations, explanations, and answers to diverse inquiries. Consequently, LLMs must be able communicate uncertainty effectively, enabling humans to appropriately calibrate their reliance on AI-generated recommendations and to understand clearly when such advice is dependable (Zhang et al., 2020; Steyvers et al., 2025; Steyvers & Kumar, 2024). Therefore, it is important to understand LLMs’ metacognitive capabilities, and to explore their capacity to communicate internal uncertainty, in order to facilitate their optimal collaboration with humans and future development potential.
Here we examine key recent findings in LLMs’ metacognitive capabilities in relation to the human literature, highlighting the methods for evaluating confidence capacities and expressions of uncertainty with emphasis on human-LLM collaboration. Throughout, we provide insights into the parallels and divergences between human and LLM metacognition, illustrate contemporary challenges in AI uncertainty quantification, and discuss potential pathways for enhancing metacognitive interactions between humans and LLMs. We end by highlighting how improving metacognitive abilities in LLMs may facilitate their future development of other capacities associated with intelligence.
2 Metacognitive Capabilities of LLMs
A key question regarding LLMs’ metacognition is whether they can accurately recognize and adequately communicate their own knowledge boundaries. Existing research is mixed in its conclusions. Some studies suggest that LLMs demonstrate limited metacognitive insight and struggle to recognize gaps in their own knowledge, leading to conclusions that LLMs lack essential metacognitive capabilities (Griot et al., 2025). Yet other findings suggest that LLMs can indeed detect their knowledge boundaries and can discriminate effectively between problems they can solve correctly and those for which they may fail (Yin et al., 2023; Kadavath et al., 2022; Steyvers et al., 2025). A contributing factor to these seemingly conflicting results is the large diversity in methods used to evaluate LLM confidence. Broadly, two approaches dominate current research: explicit confidence elicitation and implicit confidence estimation.
Explicit confidence elicitation involves prompting the model to verbalize its level of certainty—either through qualitative statements (e.g., “I’m not sure”) or quantitative confidence judgments expressed as percentages or probabilities (e.g., “I’m 70% sure”) (Cash et al., 2024; Griot et al., 2025). These outputs are generated via text, relying on the model’s ability to represent and articulate its own uncertainty in language. In contrast, implicit confidence estimation seeks to infer model confidence by measuring the likelihood assigned to tokens at the output layer of the LLM (Steyvers et al., 2025). For example, when answering a multiple-choice question with options A, B, C, and D, a confidence measure can be derived from the probability distribution over these choices. This method extends to open-ended questions through the p(true) approach (Kadavath et al., 2022), where the model first generates an answer and is then separately queried about its truthfulness. The confidence score is then calculated based on the probability that the model assigns to“true” versus “false” responses. These behaviors suggest the presence of metacognition through serving as implicit indicators, similar to how indirect evidence has been interpreted in non-human animal research: rats can “report” higher confidence in a decision by waiting longer for a food reward, and their behavioral patterns precisely map onto explicit confidence reports in humans and monkeys (Stolyarova et al., 2019). Studies consistently find that implicit confidence measures derived from token likelihoods tend to exhibit greater metacognitive sensitivity (see next section) compared to verbalized confidence elicited through explicit prompting (Xiong et al., 2024). This discrepancy highlights an important distinction between what models internally “know” (or represent) and what they explicitly express.
Additionally, metacognitive performance—by either of these approaches—is not uniform across different LLMs. Differences in model architecture and scale significantly impact the degree of apparent metacognitive capability. Larger and more sophisticated models tend to show improved calibration—meaning their expressed confidence better matches their actual accuracy—as well as heightened metacognitive sensitivity—meaning their confidence judgments can better differentiate between correct and incorrect responses (Kadavath et al., 2022; Xiong et al., 2024). These observations suggest that scaling up model size and optimizing architectural components may enhance LLMs’ metacognitive capacities. To fully comprehend and measure LLM metacognition, we must carefully think about evaluation methods and be aware of the inherent differences between models.
3 Methods to Assess Confidence-Accuracy Relationship
Several metrics have been used to assess the relationship between confidence and accuracy across both humans and AI systems. While these metrics differ across disciplines, with some metrics originating in computer science and others in cognitive science, the metrics reveal two key facets of metacognition, briefly introduced above: metacognitive calibration and metacognitive sensitivity (Fleming, 2023; Lee et al., 2024; Li & Steyvers, 2025). For visual reference, Figure 1 illustrates both concepts, and compares them to empirical results for GPT-3.5 on a multiple-choice question-answering task. In this section, we describe these facets in more detail.
Metacognitive sensitivity quantifies how “diagnostic” confidence judgments are of decisional accuracy—i.e., whether they reliably discriminate between correct or incorrect answers (Figure 1, top row). Within the human literature, metacognitive sensitivity metrics include phi ($\phi$) correlation (i.e., the correlation between accuracy and confidence across trials), the area under the type 2 receiver operating characteristic curve (AUC)—corresponding to the probability that a randomly sampled correctly answered question receives a higher confidence score than a randomly sampled incorrectly answered question—and a signal detection theoretic metric known as meta-d’ (analogous to d’ from signal detection theory), among others (Fleming & Lau, 2014). In contrast, metacognitive calibration refers to whether an observer reports a generally appropriate level of confidence given their probability of being correct. For example, if an individual—or an LLM—reports 75% confidence across multiple trials, calibration can be considered optimal when the actual proportion of correct answers in those trials is also 75% (Maniscalco et al., 2024). The Expected Calibration Error (ECE) is often used in computer science research to summarize the overall discrepancy between confidence and accuracy. ECE is typically computed by binning predictions according to confidence levels and comparing average confidence within each bin to the empirical accuracy. Calibration curves—graphs plotting model confidence against observed accuracy—are also commonly used to visualize calibration performance (Figure 1, bottom row). A perfectly calibrated system would exhibit a calibration line that falls on the diagonal (i.e., predicted confidence equals actual accuracy at all levels). Deviations from this line reflect systematic biases such as overconfidence (when predicted confidence exceeds accuracy) or underconfidence (when accuracy exceeds confidence). However, note that in the literature on human metacognition, apparent over- or under-confidence may in fact be mathematically optimal when considering reward functions or the observer’s global strategy or goals, such as whether it is more desirable to maximally avoid high-confidence errors given the consequences of such errors in the environment (Maniscalco et al., 2024).
Figure 1: Demonstrations of confidence-accuracy relationships using cartoons and an empirical example based on empirical results from GPT-3.5 from (Steyvers et al., 2025), focusing on metacognitive sensitivity and calibration. Top row: Confidence distributions for correct (green) and incorrect (blue) answers allow assessment of metacognitive sensitivity. Illustrative results show examples of different degrees of separations between the distributions reflecting different degrees of metacognitive sensitivity. The empirical results using GPT-3.5 show modest separation, with the area under the curve (AUC = 0.778) reflecting the probability that a randomly selected correct answer is assigned higher confidence than a randomly selected incorrect answer. Bottom row: Metacognitive calibration can be seen by plotting accuracy as a function of confidence. The illustrative results show examples of over-confidence, under-confidence, and properly calibrated confidence (points directly along the diagonal). The empirical results for GPT-3.5 on a multiple-choice question-answering task show overconfidence, as the model’s predicted confidence exceeds observed accuracy.
4 Comparing Human and LLM Metacognitive Architecture and Behavior
There are several notable parallels between how humans and LLMs not only generate and calibrate confidence, but also express it. These similarities may seem surprising, given the fundamental architectural and cognitive differences between humans and LLMs, yet important differences also remain; exploring these differences and their consequences on collaborative behavior may be crucial to effective human-LLM collaboration.
4.1 Similarities Between Humans and LLMs
One point of convergence may be with some mechanisms thought to generate confidence. In LLMs, one approach to estimating confidence leverages their probabilistic nature: the model can be prompted multiple times with the same question, and confidence can be inferred from the consistency of responses (Xiong et al., 2024) — similar to the other implicit measures of metacognition discussed above — or their semantic coherence (Farquhar et al., 2024). Interestingly, this approach is similar to a proposed theoretical framework for human confidence, in which subjective certainty arises from the self-consistency of internally generated candidate answers (Koriat, 2012). Although developed independently in AI and cognitive psychology, both approaches suggest that consistency across internally simulated alternatives may serve as a basis for confidence.
Another similarity concerns the outwardly-visible behavioral patterns of calibration and sensitivity. Recent work has shown that when given the same task, LLMs and humans both tend to exhibit overconfidence, and both can achieve a similar degree of metacognitive sensitivity—that is, their confidence ratings are similarly diagnostic of accuracy (Cash et al., 2024). The tendency toward overconfidence has long been observed in human cognition (Kelly & Mandel, 2024), and appears to extend to large language models as well, possibly due to inductive biases or training data characteristics (Zhou et al., 2024).
Further parallels are found in the expression and perception of linguistic uncertainty. Humans often use terms such as “likely,” “probably,” or “almost certainly” to convey probabilistic beliefs, and so do LLMs when prompted for confidence statements. Research comparing the two finds that modern LLMs match population-level human perceptions of linguistic uncertainty reasonably well when asked to translate between verbal and numeric probabilities (Belém et al., 2024).
Another important parallel is that metacognitive abilities for both humans and LLMs can be improved through training. In the case of LLMs, research has shown that confidence verbalization (specifically, metacognitive sensitivity) can be improved through finetuning approaches that reward the LLM for accurately conveying uncertainty to a listener (Stengel-Eskin et al., 2024). Similarly, humans can improve metacognitive sensitivity and calibration through training (Carpenter et al., 2019). More specifically, providing feedback, encouraging reflective reasoning, and explicitly targeting cognitive biases can reduce human miscalibration of confidence (Kelly & Mandel, 2024). However, the way in which feedback is presented may change how an observer – biological or artificial – calibrates their confidence based on the strategy the observer is trying to optimize (Maniscalco et al., 2024).
Finally, metacognition in humans is thought to rely on introspection-like processes, defined specifically by the privileged access we have to our own thoughts over those of others (i.e., the difference between metacognition and theory of mind). Similarly, it has been suggested that LLMs can better predict their own behavior than the behavior of another LLM, which some researchers interpret to imply the presence of such privileged access in the LLMs tested (Binder et al., 2024). Evidence of introspective-like capacities may also come from LLMs’ demonstrated ability to describe their own behaviors even when they are not explicitly described in their training data, including ‘backdoors’ where models show unexpected or undesirable behaviors only under certain trigger conditions (Betley et al., 2025).
4.2 Divergences Between Human and LLM Metacognition
Despite a number of parallels, there remain important differences between human and LLM metacognition. In humans, the ability to metacognitively form confidence judgments is typically thought to rest on the formation of a second-order representation: a separate evaluation or reassessment of the internal representations prompted by input information and which gave rise to a behavioral output (Peters, 2022); unless explicitly present in their architecture, LLMs may not form such second-order self-evaluative representations unless explicitly prompted to do so. Relatedly, LLMs may be less able to correctly evaluate the source of such second-order representations or judgments, suggesting they lag humans in distinguishing between metacognition and theory of mind. LLMs are prone to conflate their own beliefs with those attributed to others; that is, they are less able to separate the speaker’s belief from their own compared to humans when interpreting uncertain statements (Belém et al., 2024).
Another difference may stem from the domain generality or specificity of metacognition in humans. It is thought that some shared processes underlie metacognition about perception, memory, and cognition may exist and rely on common neural structures, while others may be domain specific—i.e., separable computational or neural modules for perceptual versus cognitive or memory metacognition (Morales et al., 2018). As LLMs are increasingly integrated into many highly different tasks and reasoning domains, attending to their domain-specific versus domain-general metacognitive capacities will become increasingly urgent (see, e.g., (Griot et al., 2025) for LLMs’ metacognitive failures in medical reasoning).
5 Communication of Uncertainty in Human-AI Interaction
To facilitate ideal collaboration between humans and LLMs, we must attend to the sources of metacognitive sensitivity and metacognitive bias in both populations—including cases where LLMs seem to engage in metacognition similarly to how humans do, but may not actually. Importantly, these behaviors and distinctions can have critical consequences for how levels of confidence can be effectively communicated between LLMs and humans.
As discussed above, metacognitive sensitivity is the degree to which confidence judgments can discriminate between right and wrong answers, which is critical to effective decision-making in humans (Fleming, 2024). For optimal interaction and humans’ trust of AI systems, LLMs thus must be able to convey to human deciders not only whether their decisions are likely to be correct (Lee et al., 2024; Steyvers et al., 2025; Li & Steyvers, 2025), but also whether they are likely to respond capably to novel tasks (Kadavath et al., 2022). Problematically, LLMs appear reluctant to express uncertainty (Zhou et al., 2024). Because humans rely heavily on linguistic uncertainty expressions (Steyvers et al., 2025; Zhou et al., 2024), the absence of expressions of uncertainty may raise humans’ reliance on model outputs even beyond the already-overconfident judgments the models express. A potential reason for LLMs’ lack of uncertainty expression may stem from training protocols, which fine-tune LLM outputs stylistically to align with human preferences for such responses—which themselves present biases against responses that contain uncertainty (Zhou et al., 2024). Unfortunately, this problem may be further exacerbated as LLMs are used for increasingly challenging applications, potentially by increasingly non-expert users. Because individuals who do not possess topical expertise are less able to correctly assess the expertise of others (Bower et al., 2024), non-expert individuals working with LLMs on complex problems might be especially susceptible to influence by superficial features of their response, like the lack of uncertainty expressions or even the length of the answer (Steyvers et al., 2025). Humans and LLMs may also rely on different sets of cues when assessing their confidence in other humans, such as humans’ reliance on the time it takes to make a response (Tullis, 2018); these cues likely will not be used in the same way by LLMs. Together, these differences in the assumed computations and inputs to metacognition may strongly impact how humans integrate LLMs’ expressed confidence into their own beliefs and decisions.
Overall, it is clear that improving AI metacognition is a key priority: LLMs must be able to differentiate correct responses from incorrect ones. Yet our research trajectory must exceed simply improving LLMs’ self-evaluation capacities if they are to effectively collaborate with humans. Imbuing LLMs with appropriate metacognitive capacities must also include directed research into their communication of uncertainty to human users, and explicit comparisons between how humans and LLMs evaluate their own uncertainty. New tasks and evaluation strategies may be beneficial in driving such development, such as building LLM capacities to recognize and name skills required to solve the task at hand (e.g., mathematical problems) (Didolkar et al., 2024). Training regimes which drive alignment between LLMs’ verbalized confidence and the perceived confidence by humans (Stengel-Eskin et al., 2024), or which emphasize LLMs’ capacities to detect questions that are beyond the scope of their knowledge base or are unanswerable (Zhang et al., 2024), may also be powerful paths forward.
6 Future Benefits of Improved AI Metacognition
Beyond the importance of improving LLMs’ metacognitive capacities to facilitate their effective integration into human-AI joint decision-making, imbuing LLMs—or any AI system—with improved metacognition may also strongly support their approach to one ultimate target: generalized intelligence. In humans, metacognitive capacities facilitate goal-directed behaviors including learning, information-seeking, and more. For example, cognitive science has long recognized the role of metacognition in driving self-directed learning, which allows us to focus effort on acquiring information that we do not yet possess—with strong parallels to active learning AI algorithms that can optimally select their own training data to maximize efficient acquisition of coherent skills or beliefs (Gureckis & Markant, 2012). Metacognitive-like evaluations of uncertainty can drive information seeking behavior even in perception and learning, allowing agents to focus on acquiring knowledge to minimize uncertainty in their internal representations of the world (Schulz et al., 2023)—a form of curiosity-driven behavior. Confidence signals can also help agents learn in reinforcement learning contexts through explicit calculation of confidence-based prediction errors (Ptasczynski et al., 2022). Finally, meta-evaluations of one’s own metacognitive abilities can also drive humans’ learning (Recht et al., 2025), and the same could be true for AI systems. It is clear that promoting LLMs’ metacognitive capacities may have profound benefit for the development of smarter, more capable AI.
7 Author Contributions
M.S. and M.A.K.P. jointly wrote this manuscript.
8 Acknowledgements
This work was partially supported by a Fellowship in the Brain, Mind, & Consciousness program from the Canadian Institute for Advanced Research. The funding agency had no role in the preparation of this manuscript.